You are here: Foswiki>PartnerObserving Web>ItalyQueue2013>ObsInaf20130929 (29 Sep 2013, EleanoraSani)Edit Attach

INAF queue observing, 2013 Sep 28-29

Observer: E. Sani, F. Cusano
Telescope Operator: D. Gonzales-Huerta
Support Astronomer: D. Thompson (on call)
Instrument: LUCI

Summary

MODS1 is currently available in red or blue grating only.
At the sunset clear sky.
At the beginning of sceintific operations the average seeing on the GCS is 0.6".
Hanging of the Telescope Control GUI. We call the SA. We figure out by specifiyng 'wait for collimation' 'OFF' in the Telescope Specific.
Condensation on the camera. We call J. Morris that have to switch on compressors. These were non-active due to the fire alarm happened in the evening.

Data Summary

Project_RUN Objects Filter ExpTime Notes
Cresci_8B Hip107734 210_zJHK, H, K   for DSF2237b
Cresci_8B DSF2237b 210_zJHK, K 30m within the specs
Cresci_8B DSF2237b 210_zJHK, H 3h within the specs. target Cresci_8B COMPLETED
Cresci_8A Hip107734 210_zJHK, K   for DSF2237a
Cresci_8A DSF2237a 210_zJHK, K 40m within the specs
Galametz_31_A1 Hip11809 200_H+K, zJspec   for CANDELS_USD18_1
Galametz_31_A1 CANDELS_USD18_1 200_H+K, zJspec 50m within the specs. target Galametz_31_A1 COMPLETED
Galametz_31_A2 CANDELS_USD18_2 200_H+K, zJspec 30m within the specs.
Galametz_31_A1 Hip11809 200_H+K, zJspec   for CANDELS_USD18_2

Details

01:09 The TO opens the dome.

01:31 acquiring skyflat Ks 20130929.0001-0016
  • from 0004-0008 note agw inside the field
  • from 0010-0016 good

02:08 slew to standard Cresci_8B
  • executeLUCI tell2B_mos.acq 0017-0019

02:17 executeLUCI tellB_mos_210_K1.spec 0020-0023
  • central wavelenght 2.17 in K
  • seeing 0.6"

02:26 executeLUCI tellB_mos_210_H2.spec 0024-0027

02:30 slewing to science CresciB
  • executeLUCI campoB_mos.acq 0028-0031

02:45 executeLUCI campoB_mos_210_K5.spec 0032-0037
  • seeing 0.7"

03:19 executeLUCI campoB_mos_210_H_3.spec 0038-0049
  • average seeing 0.6" on the GCS

04:27 executeLUCI campoB_mos_210_H_4.spec 0050-0061
  • average seeing 0.8" on the GCS

04:34 executeLUCI campoB_mos_210_H_4.spec 0062-0073

06:43 Slew to the standard for Cresci_8A 0074-0076
  • executeLUCI tellA_mos.acq
  • too far from the target

06.52 executeLUCI tell2A_mos.acq 0077-0079

06.59 executeLUCI tellA_mos_210_K.spec 0080-0083
  • We verify that the spectrum is correctly offsetted in the bluest and reddest slits.

07.09 executeLUCI campoA_mos.acq
  • collimation problems. We ask the TO to slew to a bright star and collimate
  • we are not able to preset. The Telescope Control gui is hanging. The TO try to close the connection between telescope and LUCI.
  • It doesn't work, we have to call the SA. The SA suggest to select wait for collimation "OFF" and figure it out.

07.25 executeLUCI campoA_mos.acq 0084-0086
  • we have CONDENSATION on the acquisition image. But humidity is only 69%!
  • We have to call J. Morris and slew to the Zenith to allow to improve the compressor and shrink condensation.
  • Condensation happened because the compressors where disabled due to the FIRE allarm of the early evening.
  • 0087-0088 test image to verify the presence of condensation. Condensation is gone, we can re-acquire Cresci_8A

07.54 executeLUCI campoA_mos.acq 0089-0093

08:15 executeLUCI campoA_mos_210_K_1.acq 0094-0101
  • average seeing on the GCS 0.6"

09:01 We slew to standard star for Galametz_31_A1
  • executeLUCI CANDELS_USD18_A1_tell_mos.acq 0102-0105

09:14 executeLUCI CANDELS_USD18_A1_tell_mos_zJspec.spec 0106-0108

09:18 executeLUCI CANDELS_USD18_A1_mos.acq 0109-0112

09:32 executeLUCI CANDELS_USD18_A1_mos_zJspec_2.spec 0113-0114
  • GCS gui not responding. We are only traking.
  • GCS restarted, to guide again we do have to re-acquire the target.

09:53 executeLUCI CANDELS_USD18_A1_mos.acq 0115-0117
  • we commit -0.12-2.0 and directly integrate the spectra

10:00 executeLUCI CANDELS_USD18_A1_mos_zJspec_2.spec 0118-0125
  • average seeing on the GCS 0.5"

10:46 We slew to Galametz_31_A2
  • executeLUCI CANDELS_USD18_A2_mos.acq 0126-0130

11:09 executeLUCI CANDELS_USD18_A2_mos_zJspec_2.spec 0131-0136
  • average seeing on the GCS 0.7"

11:44 Slew to standard star for Galametz_31_A2
  • executeLUCI CANDELS_USD18_A2_tell_mos.acq 0137-0140

11:56 executeLUCI CANDELS_USD18_A2_tell_zJspec.spec 0141-0143

12.00 executeLUCI EndOfNight

-- EleanoraSani - 29 Sep 2013
Topic revision: r4 - 29 Sep 2013, EleanoraSani
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback